OpenStudy (anonymous):

(x^3 - 2x^2 - 2x +1) / (x^2 + 2x + 1) The end result should be (x^2 - 3x +1) / 2, I just don't know how to get there..

OpenStudy (anonymous):

(x^3 - 2x^2 - 2x +1) / (x^2 + 2x + 1)=(x+1)(x^2 - 3x +1)/ (x+1)^2 =(x^2 - 3x +1) / x+1

OpenStudy (anonymous):

There must be a mistake in my book, at the end it says it's /2, not / x+1, but the result you gave me is the right one.. thank you! :)